OpenStudy (anonymous):

Any help is appreciated Simplify sqrt(125x^7) and (2x+24)/(3x+36)

mathslover (mathslover):

\(\sqrt{125x^7} = \sqrt{25 * 5 * x^6 * x } \) Right?

OpenStudy (anonymous):

can you help with the second one please and thank you

OpenStudy (agent0smith):

Still need help on the second?